A cubical block of wood 10cm on a side floats at the interference bet ween oil and water with its lower surface 1.50 cm below the interference. The density of the oil is 790kg/m^3. What is gauge pressure at lower surface of the block?

Grade:11

2 Answers

Arpit Kumar
10 years ago
Gauge pressure is the pressure above atmospheric pressure,mathematically,absolute pressure=atmospheric pressure + gauge pressure.height of water below is 1.5cm and height of oil above is 8.5cm According to your question absolute pressure=atmospheric pressure + pressure of oil + pressure of water=101325Pa + density of oil *g*height+density of water*g*height=790*9.8*0.085+1000*9.8*1.5=102130.07 Pa This is absolute pressure. Gauge pressure=102130.07-101325=805.07 Pa
chaitnyakishore
10 years ago
Gauge pressure is the pressure above atmospheric pressure. gauge pressure =density =(790*10*0.085)+(1000*10*0.015)=821.5Pa